# Homogeneous braids are visually prime

### v1

@@ Abstract.
We show that closures of homogeneous braids are visually prime, addressing a question of Cromwell. The key technical tool for the proof is the following criterion concerning primeness of open books, which we consider to be of independent interest. For open books of $3$–manifolds the property of having no fixed essential arcs is preserved under essential Murasugi sums with a strictly right-veering open book, if the plumbing region of the original open book veers to the left. We also provide examples of open books in $S^{3}$ demonstrating that primeness is not necessarily preserved under essential Murasugi sum, in fact not even under stabilizations a.k.a. Hopf plumbings. Furthermore, we find that trefoil plumbings need not preserve primeness. In contrast, we establish that figure-eight knot plumbings do preserve primeness.
@@ 1. Introduction
Knot theorists have a somewhat ambivalent relationship with the most common way of representing knots, i.e. by knot diagrams. The latter are the result of projecting a knot—a circle embedded in $\mathbb{R}^{3}$—to a generic $2$–dimensional plane retaining crossing information. Indeed, many simple 3–dimensional properties, e.g. primeness, are often hard to discern from a diagram. However, for particular classes of knot diagrams, one can read off a given property directly from the diagram. Concerning primeness, the two first key results are due to Menasco [Men84] and Cromwell [Cro93], respectively: they show that primeness of links—embeddings of a collection of circles in $\mathbb{R}^{3}$—represented by alternating and braid positive diagrams can be read off from such diagrams. More precisely, if the link is not prime, i.e. if there exists a decomposition sphere, then there exists a decomposition sphere that arises from a decomposition circle in the diagram. These results are condensed into the slogan “alternating links and positive braids are visually prime”. Later, Ozawa [Oza02] extended Cromwell’s result to positive diagrams. Cromwell more generally conjectured a link diagram to be visually prime if its Seifert genus is realized by a Seifert surface coming from Seifert’s algorithm. More precisely, given a link $L$ with a diagram $D$ for which Seifert’s algorithm yields a minimal Seifert surface, the following ought to hold: if $L$ is not prime, then $D$ features a decomposition circle that gives rise to a decomposition sphere [Cro93]. Cromwell points out that for diagrams arising from braids, Seifert’s algorithm yields a minimal Seifert surface if and only if the braid is homogeneous. An $n$–stranded braid word $\beta$—an expression in Artin’s standard generators $\sigma_{1},\dots,\sigma_{n-1}$ and their inverses [Art25]—is homogeneous if for each $i$ only one of $\sigma_{i}$ and $\sigma_{i}^{-1}$ occurs. We resolve Cromwell’s conjecture for diagrams that arise from braids, that is, we prove the following.
@@
Let $(\Sigma,\phi)$ be an open book—a pair consisting of an oriented compact surface $\Sigma$ with non-empty boundary, called the page, and an orientation preserving self-diffeomorphism $\phi$, called the monodromy. For the rest of the introduction and Section 2, we assume that the oriented $3$–manifold $M_{\phi}$ associated with the open book is irreducible (e.g. $M_{\phi}\cong S^{3}$) for sake of exposition. Here, the associated $3$–manifold $M_{\phi}$ is the result of building the mapping torus of $\phi$ and collapsing its boundary components to a union of circles, called the binding. The binding, understood as a link in $M_{\phi}$, is prime if the open book has no fixed essential arcs, where an arc $a$—a properly embedded closed interval—in $\Sigma$ is said to be essential if it cannot be isotoped relative boundary into $\partial\Sigma$, and is said to be fixed if $\phi(a)$ and $a$ are isotopic relative boundary.
The so inclined reader might hope that primeness is preserved under stabilization as follows. Assume $(\Sigma,\phi)$ has no fixed essential arcs and let $(\widetilde{\Sigma},\widetilde{\phi})$ be a stabilization (also known as Hopf plumbing) along some essential arc. Then, surely, $(\widetilde{\Sigma},\widetilde{\phi})$ also has no fixed essential arc; in other words its binding is again a prime link in $M_{\widetilde{\phi}}\cong M_{\phi}$?! This surmise is e.g. true if $\phi$ is positive and the stabilization is positive—that is a key input in Ito’s reproof of Cromwell’s result for positive braids [Ito22, Appendix A]; see also Remark 4.7 for how Ito’s argument relates to ours. However, it is false in general, as the following example shows.
@@
In this example we consider open books with associated $3$–manifold diffeomorphic to $S^{3}$, which we describe by their bindings: fibered links in $S^{3}$. Consider the prime fibered knot in $S^{3}$ arising as closure of the homogeneous 3–stranded braid $\sigma_{1}^{-2}\sigma_{2}\sigma_{1}^{-1}\sigma_{2}^{2}$, which is $6_{3}$ in the knot tables; see top left of Figure 1.
@@
The main technical work in our proof to establish Theorem 1.1 lies in providing a criterion that assures that indeed $(\widetilde{\Sigma},\widetilde{\phi})$ has no fixed essential arc. In fact, the criterion more generally works for Murasugi sums, a generalization of stabilization (which we recall below), and in the proof of Theorem 1.1 we exploit the natural Murasugi sum structure of the closure of homogeneous braids.
@@
As an illustration of 1.3 we obtain that so-called fibered arborescent links—links arising from iterative plumbing of Hopf bands guided by a plane tree; see [BS16, Ch. 12][Baa05]—are all prime. We see this as an illustration rather than a new result, and note that primeness, in fact, hyperbolicity is known for all but three exceptional families of arborescent knots; compare with Futer and Guéritaud’s work [FG09]. Given that 1.3 does not just work for $S^{3}$, but is stated in terms of open books with any associated $3$–manifold, we take the opportunity to introduce a notion of a tree of open books, essentially encoding the Murasugi summing of strictly right-veering and strictly left-veering open books along disjoint essential summing regions (see Section 4). We then show, using the criterion, that the resulting open book has no fixed essential arcs, and is thus, in particular, prime (see Theorem 4.5). This simultaneously generalizes the key idea in the proofs of Theorem 1.1 and of primeness of fibered arborescent links.
@@
We discussed that the property of an open book having no fixed essential arcs is not necessarily preserved by essential Hopf plumbing; see Example 1.2. It turns out that the same is true for essential trefoil plumbings, which we understand to mean the following. Starting with an open book $(\Sigma,\phi)$, plumb a positive Hopf band along an essential arc $a$ in $\Sigma$, and to the resulting open book $(\widetilde{\Sigma},\widetilde{\phi})$ plumb a second positive Hopf band along an essential arc that is disjoint from $\Sigma\subset\widetilde{\Sigma}$. As the second arc is unique up to isotopy, the resulting open book only depends on the choice of $a$, and we call it the positive trefoil plumbing of $({\Sigma},{\phi})$ along $a$.
@@
In other words, there exists a prime fibered knot in a $3$–manifold $M_{\phi}$ that features an essential trefoil plumbing for which the resulting knot in $M_{\phi}$ is not prime.
Similarly, given an open book $(\Sigma,\phi)$ and an arc $a$ in $\Sigma$, we call the same construction as above, but replacing the second positive Hopf band by a negative Hopf band, the figure-eight plumbing of $({\Sigma},{\phi})$ along $a$. In contrast to trefoil plumbings, figure-eight plumbings preserves the property of having no fixed essential arcs, as a rather immediate application of 1.3 shows.
@@
The authors were surprised by this dichotomy between trefoil plumbings and figure-eight plumbings. However, we do not know whether an open book as in Proposition 1.4 can be found such that the associated $3$–manifold is $S^{3}$.
@@
In Section 2, we sketch a proof of Theorem 1.1 using 1.3. In Section 3, we give detailed definitions concerning open books, Murasugi sums, veeringness, and prove 1.3. In Section 4, we describe an iterative construction of Murasugi summing guided by a tree of open books, and we establish Theorem 4.5, which states that the resulting open book is prime, whenever all involved open books are strictly veering and and the summing regions are essential. As an application, we discuss primeness of arborescent fibered links and prove Proposition 1.5. In Section 5, we give an introduction to homogeneous braids and decomposition circles. Then we prove several lemmas on the Murasugi sum structure of homogeneous braids, which we use to give a detailed proof of Theorem 1.1 by inductively showing that prime diagrams of homogeneous braids are obtained by repeated Murasugi sums satisfying the criterion. We do so using the language of trees of open books developed in Section 4. Finally, in Section 6, we prove Proposition 1.4.
@@ Acknowledgments
PF and MOR gratefully acknowledge financial support by the SNSF Grant 181199.
@@ 2. Outline of the proof of visual primeness of homogeneous braids
We provide a sketch of the proof of Theorem 1.1, relying on what we hope the reader will find instructive examples, namely the $5$–stranded homogeneous braids
@@
and invoking 1.3. In this sketch we readily switch between open books $(\Sigma,\phi)$ for which the associated $3$–manifolds is diffeomorphic to $S^{3}$ and their associated fiber surfaces in $S^{3}$.
@@
Firstly, $\Sigma(\beta)$ is the iterative Murasugi summing of $n-1$ open books with pages that are the fiber surfaces $\Sigma(2,k_{1}),\dots\Sigma(2,k_{n-1})$ of the torus links $T(2,k_{i})$, where $k_{i}\in\mathbb{Z}\setminus\{-1,0,1\}$. In the case of our two example braids, we find $k_{1}=k_{4}=-3$, $k_{2}=3$, and $k_{3}=5$; see Figure 3.
@@
Secondly, we describe a summing procedure that uses the same summing regions, but first iteratively sums together $\Sigma(2,k_{i})$, $\Sigma(2,k_{i+1})$, $\Sigma(2,k_{i+l})$, where $i$ is chosen minimally and $l$ is chosen maximally such that $k_{i},\dots,k_{i+l}$ have the same sign, yielding open books given by fiber surfaces $\Sigma_{1},\dots,\Sigma_{m}$ for some $m\leq n-1$. If the summing regions that are used define essential Murasugi sums, then these open books are all either strictly right-veering or strictly left-veering since they are iterative essential Murasugi sums of open books that all are strictly right-veering or all are strictly left-veering; see Lemma 4.6. In fact, $\Sigma_{1},\dots\Sigma_{m}$ alternate between being strictly right-veering and strictly left-veering. In the case of $\beta_{\mathrm{prime}}$, this yields the $3$ fiber surfaces $\Sigma_{1}$, $\Sigma_{2}$, and $\Sigma_{3}$ depicted in Figure 4.
@@ 3.1. Preliminaries on open books, fixed arcs and primeness
We call an open book $(\Sigma,\phi)$ prime if there does not exist an separating essential arc that is fixed by $\phi$. Equivalently, $(\Sigma,\phi)$ is prime if there does not exist an embedded $2$–sphere in the $3$–manifold associated with the open book that transversely intersects the page in a separating essential arc. Here, the $3$manifold $M_{\phi}$ associated with the open book $(\Sigma,\phi)$ is the closed oriented $3$–manifold resulting as a quotient of the mapping torus of $\phi$ as follows: collapse the boundary tori of the mapping torus $\Sigma\times[0,1]/(x,0)\sim(\phi(x),1)$ to circles—called the binding—by collapsing the $[0,1]$–direction to a point. The page is considered as a surface in $M_{\phi}$, aka a fiber surface for the binding, via the embedding of $\Sigma$ induced by $x\mapsto(x,0)$. In fact, a fixed essential arc is separating exactly if the corresponding $2$–sphere separates the $3$–manifold. In this text we focus on the question of whether an open book has no fixed essential arcs, as this is what we can naturally track from the point of view of open books, rather than separating arcs. While this is a priori a stronger condition than primeness, note that in case the resulting $3$–manifold $M_{\phi}$ does not feature an embedding of a $2$–sphere that is non-separating, which is equivalent to $M$ not having a $S^{1}\times S^{2}$ summand in its prime decomposition, the two notions agree.
Now, let $L$ be a link in an oriented closed $3$–manifold $M$. We say $L$ is fibered if its complement fibers over the circle, that is, there exists a fibration $\pi\colon M\setminus L\rightarrow S^{1}$ where the fibers $F_{\theta}=\pi^{-1}(\theta)$ are the interior of diffeomorphic compact surfaces with common boundary $L$. Equivalently, there exists an open book such that $M_{\phi}$ is diffeomorphic to $M$ by an orientation preserving diffeomorphism that orientation preservingly maps the binding to $L$. In this case, the page can be taken to be a surface whose interior is diffeomorphic to the fiber $F_{\theta}$, and the monodromy is the (equivalence class of the) return map of the fibration. Here, equivalence of open books is the equivalence relation induced by boundary preserving isotopy and conjugation: open books $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ are conjugate if there exist an orientation preserving diffeomorphism $f\colon\Sigma_{1}\to\Sigma_{2}$ such that $\phi_{2}\circ f=f\circ\phi_{1}$. Associating to an open book $(\Sigma,\phi)$ the $3$–manifold $M_{\phi}$ and its binding induces a bijection between open books up to equivalence and fibered links $L$ in $3$–manifolds $M$ up to orientation preserving diffeomorphisms of pairs $(M,L)$. We note that a fibered link in $S^{3}$ is prime (as defined in knot theory; compare Section 5) if and only if the corresponding open book is prime, justifying the above definition of primeness for open books.
@@ 3.2. Murasugi sums and representing fixed essential arcs therein
Originally, the Murasugi sum was understood as an operation on surfaces embedded in $S^{3}$ [Mur58, Mur63]. However, as Gabai proved that it preserves fiberedness [Gab83], one can define it for open books of $S^{3}$, and the definition extends to open books without restriction on the associated $3$–manifolds.
@@
Let $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ be two open books. Let $n\geq 1$ be an integer and $P$ be a $2n$–gon whose boundary has sides $s_{1},\dots,s_{2n}$. For $i=1,2$, let $f_{i}$ be an embedding of $P$ into $\Sigma_{i}$ such that, for $j=1,\dots,n$, $f_{i}(s_{2j})\subset\partial\Sigma_{i}$, and $f_{i}(s_{2j-1})$ is an arc properly embedded in $\Sigma_{i}$. We identify $P$ with its image in $\Sigma_{1}$, $\Sigma_{2}$, and $\Sigma$, and often suppress $f_{i}$ in the notation. Let $\Sigma$ be the result of gluing $\Sigma_{2}$ to $\Sigma_{1}$ along $P$, i.e. $\Sigma\coloneqq\Sigma_{1}\cup_{f_{1}\circ f_{2}^{-1}\colon f_{2}(P)\to f_{1}(P)}\Sigma_{2}$, and set $\phi\coloneqq\phi_{2}\circ\phi_{1}$. Here, in slight abuse of notation, we write $\phi_{1}$ to be the diffeomorphism of $\Sigma$ that is obtained by doing $\phi_{1}$ on $\Sigma_{1}$ and the identity on $\Sigma_{2}$ (and similarly for $\phi_{2}$). The open book $(\Sigma,\phi)$ is called the Murasugi sum of $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ along $P$. Moreover, if for $i=1,2$, $j=1,\dots,n$, $f_{i}(s_{2j-1})$ is essential in $\Sigma_{i}$, we say the Murasugi sum and the summing region are essential. In another abuse of notation, we denote the sides of $P\subset\Sigma$ by $s_{i}$. Using this notation, a Murasugi sum respectively the plumbing region $P$ is essential, if all the sides $s_{i}$ of $P\subset\Sigma$ are essential in $\Sigma$.
If an $n$–Murasugi sum (i.e. a Murasugi sum along a $2n$–gon) with $n\geq 2$ is not essential, then it is equivalent to an $(n-1)$–Murasugi sum. Note that a $1$–Murasugi sum is a boundary connected sum. In this case, if neither of the summands is a disk, we can always find a fixed essential arc (the one that separates the two surfaces). Thus we can restrict ourselves to essential Murasugi sums.
@@
Under the correspondence between open books and fibered links in $3$–manifolds, the Murasugi sum of $L_{1}$ in $M_{1}$ with $L_{2}$ in $M_{2}$ amounts to the $3$–manifold $M$ given as the connected sum of $M_{1}$ and $M_{2}$ and the link $L$ in $M$ defined as follows [Gab83]. Choose pages in $M_{1}$ and $M_{2}$ and closed balls $B_{1}$ and $B_{2}$ in $M_{1}$ and $M_{2}$, respectively, such that $B_{i}$ intersects the page exactly in the summing region $P$. Writing $M_{i}^{\circ}\coloneqq M_{i}\setminus\mathrm{int}(B_{i})$, we take $M$ to be $M_{1}\cup_{F}M_{2}$, where $F$ is an orientation reversing diffeomorphism identifying the boundaries of $B_{i}$ such that it extends the identifications of $\partial P$ viewed in the two pages respectively. Then the two-sphere $S$ in $M$ defined as the quotient of $\partial B_{i}$ is the decomposition sphere for the connected sum and the link $L$ is taken to be (a smoothing of) the result of taking the quotient of $L_{1}\setminus\mathrm{int}(B_{1})\cup L_{2}\setminus\mathrm{int}(B_{1})\subset M_{1}\sqcup M_{2}$ in $M$. See Figure 5 for a visualization.
Figure 5. Model of $\Sigma$ in a ball-neighborhood of $P$ in the $3$–manifold associated to a $3$–Murasugi sum. The $6$–gon $P$ along which we are summing is shaded in gray, and lies on a plane $E$. The page $\Sigma_{1}$ lies below $E$, and above it, lies the page $\Sigma_{2}$. In the notation of Remark 3.3, in case $M_{1}\cong M_{2}\cong S^{3}$, up to diffeomorphism $M$ can be identified with the one point compactification of $\mathbb{R}^{3}$ such that $S^{2}$ is the one-point compactification of $E$, $M_{1}\setminus\mathrm{int}(B_{1})$ and $M_{2}\setminus\mathrm{int}(B_{2})$ are the lower and upper half-spaces with boundary $E$, respectively, and the pages $\Sigma_{i}\subset M$ are surfaces in the said half-spaces.

### v2

@@ Abstract.
We show that closures of homogeneous braids are visually prime, addressing a question of Cromwell. The key technical tool for the proof is the following criterion concerning primeness of open books, which we consider to be of independent interest. For open books of 33–manifolds the property of having no fixed essential arcs is preserved under essential Murasugi sums with a strictly right-veering open book, if the plumbing region of the original open book veers to the left. We also provide examples of open books in $S^{3}$ demonstrating that primeness is not necessarily preserved under essential Murasugi sum, in fact not even under stabilizations a.k.a. Hopf plumbings. Furthermore, we find that trefoil plumbings need not preserve primeness. In contrast, we establish that figure-eight knot plumbings do preserve primeness.
@@ 1. Introduction
Knot theorists have a somewhat ambivalent relationship with the most common way of representing knots, i.e. by knot diagrams. The latter are the result of projecting a knot—a circle embedded in $\mathbb{R}^{3}$—to a generic 22–dimensional plane retaining crossing information. Indeed, many simple 3–dimensional properties, e.g. primeness, are often hard to discern from a diagram. However, for particular classes of knot diagrams, one can read off a given property directly from the diagram. Concerning primeness, the two first key results are due to Menasco [Men84] and Cromwell [Cro93], respectively: they show that primeness of links—embeddings of a collection of circles in $\mathbb{R}^{3}$—represented by alternating and braid positive diagrams can be read off from such diagrams. More precisely, if the link is not prime, i.e. if there exists a decomposition sphere, then there exists a decomposition sphere that arises from a decomposition circle in the diagram. These results are condensed into the slogan “alternating links and positive braids are visually prime”. Later, Ozawa [Oza02] extended Cromwell’s result to positive diagrams. Cromwell more generally conjectured a link diagram to be visually prime if its Seifert genus is realized by a Seifert surface coming from Seifert’s algorithm. More precisely, given a link $L$ with a diagram $D$ for which Seifert’s algorithm yields a minimal Seifert surface, the following ought to hold: if $L$ is not prime, then $D$ features a decomposition circle that gives rise to a decomposition sphere [Cro93]. Cromwell points out that for diagrams arising from braids, Seifert’s algorithm yields a minimal Seifert surface if and only if the braid is homogeneous. An nn–stranded braid word $\beta$—an expression in Artin’s standard generators $\sigma_{1},\dots,\sigma_{n-1}$ and their inverses [Art25]—is homogeneous if for each $i$ only one of $\sigma_{i}$ and $\sigma_{i}^{-1}$ occurs. We resolve Cromwell’s conjecture in the positive for diagrams that arise from braids [^1] We note that Stoimenow found a counterexample to the full conjecture [Sto08, Figure 15]., that is, we prove the following.
@@
Let $(\Sigma,\phi)$ be an open book—a pair consisting of an oriented compact surface $\Sigma$ with non-empty boundary, called the page, and an orientation preserving self-diffeomorphism $\phi$, called the monodromy. For the rest of the introduction and Section 2, we assume that the oriented 33–manifold $M_{\phi}$ associated with the open book is irreducible (e.g. $M_{\phi}\cong S^{3}$) for sake of exposition. Here, the associated 33–manifold $M_{\phi}$ is the result of building the mapping torus of $\phi$ and collapsing its boundary components to a union of circles, called the binding. The binding, understood as a link in $M_{\phi}$, is prime if the open book has no fixed essential arcs, where an arc aa—a properly embedded closed interval—in $\Sigma$ is said to be essential if it cannot be isotoped relative boundary into $\partial\Sigma$, and is said to be fixed if $\phi(a)$ and $a$ are isotopic relative boundary.
[…]
@@
In this example we consider open books with associated 33–manifold diffeomorphic to $S^{3}$, which we describe by their bindings: fibered links in $S^{3}$. Consider the prime fibered knot in $S^{3}$ arising as closure of the homogeneous 3–stranded braid $\sigma_{1}^{-2}\sigma_{2}\sigma_{1}^{-1}\sigma_{2}^{2}$, which is $6_{3}$ in the knot tables; see top left of Figure 1.
@@
[…]
@@
As an illustration of 1.3 we obtain that so-called fibered arborescent links—links arising from iterative plumbing of Hopf bands guided by a plane tree; see [BS16, Ch. 12][Baa05]—are all prime. We see this as an illustration rather than a new result, and note that primeness, in fact, hyperbolicity is known for all but three exceptional families of arborescent knots; compare with Futer and Guéritaud’s work [FG09]. Given that 1.3 does not just work for $S^{3}$, but is stated in terms of open books with any associated 33–manifold, we take the opportunity to introduce a notion of a tree of open books, essentially encoding the Murasugi summing of strictly right-veering and strictly left-veering open books along disjoint essential summing regions (see Section 4). We then show, using the criterion, that the resulting open book has no fixed essential arcs, and is thus, in particular, prime (see Theorem 4.5). This simultaneously generalizes the key idea in the proofs of Theorem 1.1 and of primeness of fibered arborescent links.
@@
[…]
@@
In other words, there exists a prime fibered knot in a 33–manifold $M_{\phi}$ that features an essential trefoil plumbing for which the resulting knot in $M_{\phi}$ is not prime.
Similarly, given an open book $(\Sigma,\phi)$ and an arc $a$ in $\Sigma$, we call the same construction as above, but replacing the second positive Hopf band by a negative Hopf band, the figure-eight plumbing of $({\Sigma},{\phi})$ along $a$. In contrast to trefoil plumbings, figure-eight plumbings preserve the property of having no fixed essential arcs, as a rather immediate application of 1.3 shows.
@@
The authors were surprised by this dichotomy between trefoil plumbings and figure-eight plumbings. However, we do not know whether an open book as in Proposition 1.4 can be found such that the associated 33–manifold is $S^{3}$.
@@
In Section 2, we sketch a proof of Theorem 1.1 using 1.3. In Section 3, we give detailed definitions concerning open books, Murasugi sums, veeringness, and prove 1.3. In Section 4, we describe an iterative construction of Murasugi summing guided by a tree of open books, and we establish Theorem 4.5, which states that the resulting open book is prime, whenever all involved open books are strictly veering and the summing regions are essential. As an application, we discuss primeness of arborescent fibered links and prove Proposition 1.5. In Section 5, we give an introduction to homogeneous braids and decomposition circles. Then we prove several lemmas on the Murasugi sum structure of homogeneous braids, which we use to give a detailed proof of Theorem 1.1 by inductively showing that prime diagrams of homogeneous braids are obtained by repeated Murasugi sums satisfying the criterion. We do so using the language of trees of open books developed in Section 4. Finally, in Section 6, we prove Proposition 1.4.
@@ Acknowledgments
We would like to thank Mark Pencovitch for pointing out Stoimenow’s example. PF and MOR gratefully acknowledge financial support by the SNSF Grant 181199.
@@ 2. Outline of the proof of visual primeness of homogeneous braids
We provide a sketch of the proof of Theorem 1.1, relying on what we hope the reader will find instructive examples, namely the 55–stranded homogeneous braids
@@
and invoking 1.3. In this sketch we readily switch between open books $(\Sigma,\phi)$ for which the associated 33–manifolds are diffeomorphic to $S^{3}$ and their associated fiber surfaces in $S^{3}$.
@@
Firstly, $\Sigma(\beta)$ is the iterative Murasugi summing of n1n-1 open books with pages that are the fiber surfaces $\Sigma(2,k_{1}),\dots\Sigma(2,k_{n-1})$ of the torus links $T(2,k_{i})$, where $k_{i}\in\mathbb{Z}\setminus\{-1,0,1\}$. In the case of our two example braids, we find $k_{1}=k_{4}=-3$, $k_{2}=3$, and $k_{3}=5$; see Figure 3.
@@
Secondly, we describe a summing procedure that uses the same summing regions, but first iteratively sums together $\Sigma(2,k_{i})$, $\Sigma(2,k_{i+1})$$\dots,$ $\Sigma(2,k_{i+l})$, where $i$ is chosen minimally and $l$ is chosen maximally such that $k_{i},\dots,k_{i+l}$ have the same sign, yielding open books given by fiber surfaces $\Sigma_{1},\dots,\Sigma_{m}$ for some $m\leq n-1$. If the summing regions that are used define essential Murasugi sums, then these open books are all either strictly right-veering or strictly left-veering since they are iterative essential Murasugi sums of open books that all are strictly right-veering or all are strictly left-veering; see Lemma 4.6. In fact, $\Sigma_{1},\dots\Sigma_{m}$ alternate between being strictly right-veering and strictly left-veering. In the case of $\beta_{\mathrm{prime}}$, this yields the $3$ fiber surfaces $\Sigma_{1}$, $\Sigma_{2}$, and $\Sigma_{3}$ depicted in Figure 4.
@@ 3.1. Preliminaries on open books, fixed arcs and primeness
We call an open book $(\Sigma,\phi)$ prime if there does not exist an separating essential arc that is fixed by $\phi$. Equivalently, $(\Sigma,\phi)$ is prime if there does not exist an embedded 22–sphere in the 33–manifold associated with the open book that transversely intersects the page in a separating essential arc. Here, the 33manifold $M_{\phi}$ associated with the open book $(\Sigma,\phi)$ is the closed oriented 33–manifold resulting as a quotient of the mapping torus of $\phi$ as follows: collapse the boundary tori of the mapping torus $\Sigma\times[0,1]/(x,0)\sim(\phi(x),1)$ to circles—called the binding—by collapsing the [0,1][0,1]–direction to a point. The page is considered as a surface in $M_{\phi}$, a.k.a. a fiber surface for the binding, via the embedding of $\Sigma$ induced by $x\mapsto(x,0)$. In fact, a fixed essential arc is separating exactly if the corresponding 22–sphere separates the 33–manifold. In this text we focus on the question of whether an open book has no fixed essential arcs, as this is what we can naturally track from the point of view of open books, rather than separating arcs. While this is a priori a stronger condition than primeness, note that in case the resulting 33–manifold $M_{\phi}$ does not feature an embedding of a 22–sphere that is non-separating, which is equivalent to $M$ not having a $S^{1}\times S^{2}$ summand in its prime decomposition, the two notions agree.
Now, let $L$ be a link in an oriented closed 33–manifold $M$. We say $L$ is fibered if its complement fibers over the circle, that is, there exists a fibration $\pi\colon M\setminus L\rightarrow S^{1}$ where the fibers $F_{\theta}=\pi^{-1}(\theta)$ are the interior of diffeomorphic compact surfaces with common boundary $L$. Equivalently, there exists an open book such that $M_{\phi}$ is diffeomorphic to $M$ by an orientation preserving diffeomorphism that orientation preservingly maps the binding to $L$. In this case, the page can be taken to be a surface whose interior is diffeomorphic to the fiber $F_{\theta}$, and the monodromy is the (equivalence class of the) return map of the fibration. Here, equivalence of open books is the equivalence relation induced by boundary preserving isotopy and conjugation: open books $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ are conjugate if there exist an orientation preserving diffeomorphism $f\colon\Sigma_{1}\to\Sigma_{2}$ such that $\phi_{2}\circ f=f\circ\phi_{1}$. Associating to an open book $(\Sigma,\phi)$ the 33–manifold $M_{\phi}$ and its binding induces a bijection between open books up to equivalence and fibered links $L$ in 33–manifolds $M$ up to orientation preserving diffeomorphisms of pairs $(M,L)$. We note that a fibered link in $S^{3}$ is prime (as defined in knot theory; compare Section 5) if and only if the corresponding open book is prime, justifying the above definition of primeness for open books.
@@ 3.2. Murasugi sums and representing fixed essential arcs therein
Originally, the Murasugi sum was understood as an operation on surfaces embedded in $S^{3}$ [Mur58, Mur63]. However, as Gabai proved that it preserves fiberedness [Gab83], one can define it for open books of $S^{3}$, and the definition extends to open books without restriction on the associated 33–manifolds.
@@
Let $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ be two open books. Let $n\geq 1$ be an integer and $P$ be a 2n2n–gon whose boundary has sides $s_{1},\dots,s_{2n}$. For $i=1,2$, let $f_{i}$ be an embedding of $P$ into $\Sigma_{i}$ such that, for $j=1,\dots,n$, $f_{1}(s_{2j-1})\subset\partial\Sigma_{1}$, $f_{2}(s_{2j})\subset\partial\Sigma_{2}$, and $f_{1}(s_{2j})\subset\Sigma_{1}$ and $f_{2}(s_{2j-1})\subset\Sigma_{2}$ are properly embedded arcs in the respective surfaces. We identify $P$ with its image in $\Sigma_{1}$, $\Sigma_{2}$, and $\Sigma$, and often suppress $f_{i}$ in the notation. Let $\Sigma$ be the result of gluing $\Sigma_{2}$ to $\Sigma_{1}$ along $P$, i.e. $\Sigma\coloneqq\Sigma_{1}\cup_{f_{1}\circ f_{2}^{-1}\colon f_{2}(P)\to f_{1}(P)}\Sigma_{2}$, and set $\phi\coloneqq\phi_{2}\circ\phi_{1}$. Here, in slight abuse of notation, we write $\phi_{1}$ to be the diffeomorphism of $\Sigma$ that is obtained by doing $\phi_{1}$ on $\Sigma_{1}$ and the identity on $\Sigma_{2}$ (and similarly for $\phi_{2}$). The open book $(\Sigma,\phi)$ is called the Murasugi sum of $(\Sigma_{1},\phi_{1})$ and $(\Sigma_{2},\phi_{2})$ along $P$. Moreover, if for $j=1,\dots,n$, $f_{1}(s_{2j})\subset\Sigma_{1}$ and $f_{2}(s_{2j-1})\subset\Sigma_{2}$ are essential arcs, we say the Murasugi sum and the summing region are essential. In another abuse of notation, we denote the sides of $P\subset\Sigma$ by $s_{i}$. Using this notation, a Murasugi sum respectively the plumbing region $P$ is essential, if all the sides $s_{i}$ of $P\subset\Sigma$ are essential in $\Sigma$.
If an nn–Murasugi sum (i.e. a Murasugi sum along a 2n2n–gon) with $n\geq 2$ is not essential, then it is equivalent to an (n1)(n-1)–Murasugi sum. Note that a 11–Murasugi sum is a boundary connected sum. In this case, if neither of the summands is a disk, we can always find a fixed essential arc (the one that separates the two surfaces). Thus we can restrict ourselves to essential Murasugi sums.
@@
Under the correspondence between open books and fibered links in 33–manifolds, the Murasugi sum of $L_{1}$ in $M_{1}$ with $L_{2}$ in $M_{2}$ amounts to the 33–manifold $M$ given as the connected sum of $M_{1}$ and $M_{2}$ and the link $L$ in $M$ defined as follows [Gab83]. Choose pages in $M_{1}$ and $M_{2}$ and closed balls $B_{1}$ and $B_{2}$ in $M_{1}$ and $M_{2}$, respectively, such that $B_{i}$ intersects the page exactly in the summing region $P$. Writing $M_{i}^{\circ}\coloneqq M_{i}\setminus\mathrm{int}(B_{i})$, we take $M$ to be $M_{1}\cup_{F}M_{2}$, where $F$ is an orientation reversing diffeomorphism identifying the boundaries of $B_{i}$ such that it extends the identifications of $\partial P$ viewed in the two pages respectively. Then the two-sphere $S$ in $M$ defined as the quotient of $\partial B_{i}$ is the decomposition sphere for the connected sum of $M$ and the link $L$ is taken to be (a smoothing of) the result of taking the quotient of $L_{1}\setminus\mathrm{int}(B_{1})\cup L_{2}\setminus\mathrm{int}(B_{1})\subset M_{1}\sqcup M_{2}$ in $M$. See Figure 5 for a visualization.
Figure 5. Model of $\Sigma$ in a ball-neighborhood of $P$ in the 33–manifold associated to a 33–Murasugi sum. The 66–gon $P$ along which we are summing is shaded in gray, and lies on a plane $E$. The page $\Sigma_{1}$ lies below $E$, and above it, lies the page $\Sigma_{2}$. In the notation of Remark 3.3, in case $M_{1}\cong M_{2}\cong S^{3}$, up to diffeomorphism $M$ can be identified with the one point compactification of $\mathbb{R}^{3}$ such that $S^{2}$ is the one-point compactification of $E$, $M_{1}\setminus\mathrm{int}(B_{1})$ and $M_{2}\setminus\mathrm{int}(B_{2})$ are the lower and upper half-spaces with boundary $E$, respectively, and the pages $\Sigma_{i}\subset M$ are surfaces in the said half-spaces.